# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v2

@@ III-B Convergence Analysis
Based on Lemma 1 in [32] and the update law of $\lambda_{i}^{k}$ in (18), it can be obtained that for any $\lambda\in\mathbb{R}_{+}^{Np}$,
$\sum_{i=1}^{M}\left\|\lambda_{i}^{k+1}-\lambda\right\|^{2}$ (23)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}+\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})-\lambda\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}\!+\!\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\lambda_{j}^{k}+\zeta_{j}^{k}-\lambda_{i}^{k})\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})-\lambda\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda+\chi^{k}\xi_{i}^{k}+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left(\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right\|^{2}+\left\|\chi^{k}\xi_{i}^{k}+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right\|^{2}\right.$
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right)^{\top}\left(\chi^{k}\xi_{i}^{k}\right)\right.$
$\qquad\left.+2\left(\bar{\lambda}^{k}-\lambda\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right.$
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\bar{\lambda}^{k}\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right),$
where $w_{ij}^{k}$ and $\xi_{i}^{k}$ are defined as
$w_{ii}^{k}:=1+\chi^{k}L_{ii},\quad w_{ij}^{k}:=\chi^{k}L_{ij},\quad\xi_{i}^{k}:=\sum_{j\in\mathcal{N}_{i}}L_{ij}\zeta_{j}^{k}.$ (24)
@@
Using (25), it can be derived that
$\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right\|^{2}$ $=\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\left(\lambda_{j}^{k}-\lambda\right)\right\|^{2}$ (27)
$\leq\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda\right\|^{2}.$
@@
$\qquad\left.+J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})\right)$
$\leq$ $\gamma^{k}\sum_{i=1}^{M}\left(\left(\bar{\lambda}^{k}-\lambda_{i}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})+\left(\lambda_{i}^{k}-\bar{\lambda}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right.$
$\qquad\left.+J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\left(\bar{\lambda}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right.$
$\qquad\left.-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})-\lambda^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)$
@@
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda\right\|^{2}|\mathcal{F}^{k}\right]$ (29)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda\right\|^{2}\!+d^{k}\!+2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda)\right),$
where $d^{k}$ is given by
$d^{k}=$ $(\chi^{k})^{2}\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}+M(\gamma^{k})^{2}C_{g}^{2}$ (30)
$+6C_{g}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|.$
Based on Assumption 4, Theorem 1, and the conditions for $\chi^{k}$ and $\gamma^{k}$ in (21), it can be concluded that $d^{k}$ is summable, i.e., $\sum_{k=0}^{\infty}d^{k}<\infty$.
Plugging the optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ into (29) and utilizing the Saddle-Point Theorem (12), we can arrive at
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda^{*}\right\|^{2}|\mathcal{F}^{k}\right]$ (31)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda^{*}\right\|^{2}\!+\!d^{k}\!+\!2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right),$
and
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda^{*}\right\|^{2}|\mathcal{F}^{k}\right]$ (32)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda^{*}\right\|^{2}\!+\!d^{k}\!+\!2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\!-\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right).$
According to Lemma 1, (31), and (32), it can be concluded that the following relationships hold almost surely:
$\sum_{i=1}^{M}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right)<\infty,$ (33)
$\sum_{i=1}^{M}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right)<\infty.$
Since $\gamma^{k}$ is non-summable, we have that ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})$ converge to zero almost surely. ∎
@@
The conditions for the weakening factor sequence $\{\chi^{k}\}$ and the step-size sequence $\{\gamma^{k}\}$ in Theorems 1 and 2 can be satisfied, e.g., by selecting $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\gamma^{k}=\frac{c_{4}}{1+c_{5}k}$ with any $c_{1}>0$, $c_{2}>0$, $10.5<c_{3}<$ 1, $c_{4}>0$, and $c_{5}>0$. Note that the design of $\chi^{k}$ in this example is identical to the one in Remark 1. Therefore, the sequences $\{\chi^{k}\}$, $\{\gamma^{k}\}$, and $\{\nu^{k}\}$ can be meticulously tailored to meet all conditions required by Assumption 4 and Theorems 1, 2.
@@ III-C Privacy Analysis
Based on the adjacency concept delineated in Definition 1, we can establish two adjacent distributed optimization problems, denoted as $\mathcal{P}$ and $\mathcal{P}^{\prime}$. There is only one signal that differs between these two problems, and without loss of generality we denote it as $g_{i}$ in $\mathcal{P}$ and $g^{\prime}_{i}$ in $\mathcal{P}^{\prime}$. According to the third condition of Definition 1, signals $g_{i}$ and $g^{{}^{\prime}}_{i}$ are required to exhibit similar behaviors around the optimal solution, i.e., $g_{i}$ and $g^{{}^{\prime}}_{i}$ should converge to each other if the algorithm can ensure convergence to the optimal solution. Therefore, according to the proven convergence in Theorem 2, we can formalize this condition by stipulating the existence of a constant $C>0$ such that
$\|g_{i}(\tilde{\bm{u}}^{k}_{i})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k}_{i})}\|_{1}\leq C\chi^{k}$ (34)
@@
For Algorithm 3, an execution is represented as $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\ldots\}$ with $\vartheta^{k}=\lambda^{k}=\begin{bmatrix}(\lambda_{1}^{k})^{\top},\cdots,(\lambda_{M}^{k})^{\top}\end{bmatrix}^{\top}$. An observation sequence is denoted as $\mathcal{O}=\{o^{0},o^{1},\ldots\}$ with $o^{k}=\hat{\lambda}^{k}=\begin{bmatrix}(\hat{\lambda}_{1}^{k})^{\top},\cdots,(\hat{\lambda}_{M}^{k})^{\top}\end{bmatrix}^{\top}$ (note that $\hat{\lambda}_{i}^{k}=\lambda_{i}^{k}+\zeta_{i}^{k}$, as detailed in Algorithm 3). Similar to the sensitivity metric proposed for constraint-free distributed optimization in [22], we formulate the sensitivity of Algorithm 3 in the following manner:
@@
$\Delta^{k}:=\sup\limits_{\mathcal{O}\in\mathbb{O}}\left\{\sup\limits_{\vartheta\in\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O}),\>\vartheta^{{}^{\prime}}\in\mathcal{R}_{\mathcal{P}^{{}^{\prime}},\vartheta^{0}}^{-1}(\mathcal{O})}\|\vartheta^{k}-\vartheta^{\prime k}\|_{1}\right\},$ (35)
@@
In Algorithm 3, at each iteration $k$, if each subsystem’s DP noise vector $\zeta_{i}^{k}\in\mathbb{R}^{Np}$ comprises $Np$ independent Laplace noises with parameter $\nu^{k}$, satisfying $\sum_{k=1}^{T_{0}}\frac{\Delta^{k}}{\nu^{k}}\leq\bar{\epsilon}$ for some $\bar{\epsilon}>0$, then Algorithm 3 achieves $\epsilon$-differential privacy with the cumulative privacy level for iterations $0\leq k\leq T_{0}$ less than $\bar{\epsilon}$.
@@
The proof of this lemma follows the same reasoning as that of Lemma 2 in [22]. ∎
@@
(Lemma 4, [33]) Let $\{\psi^{k}\}$ be a non-negative sequence, and $\{a^{k}\}$ and $\{b^{k}\}$ be positive sequences satisfying $\sum_{k=0}^{\infty}a^{k}=\infty$, $\lim_{k\rightarrow\infty}a^{k}=0$, and $\frac{b^{k}}{a^{k}}$ converges to zero with a polynomial rate. If there exists a $\bar{K}\geq 0$ such that $\psi^{k+1}\leq(1-a^{k})\psi^{k}+b^{k}$ holds for all $k\geq\bar{K}$, then it follows that $\psi^{k}\leq\bar{C}\frac{b^{k}}{a^{k}}$ for all $k$, with $\bar{C}$ being some constant.
@@
Suppose the conditions of Theorem 1 hold. If every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$ satisfies Assumption 4, then the following results hold:
  1. 1.
    For any finite number of iterations $T$, Algorithm 3 ensures $\epsilon$-differential privacy, and the cumulative privacy budget is bounded by $\epsilon\leq\sum_{k=1}^{T}\frac{C\varsigma^{k}}{\nu^{k}}$, where $\varsigma^{k}:=\sum_{s=1}^{k-1}\Pi_{q=s}^{k-1}(1-\chi^{q}\bar{L})\gamma^{s-1}\chi^{s-1}+\gamma^{k-1}\chi^{k-1}$, $\bar{L}:=\min_{i\in\mathbb{Z}_{1}^{M}}|L_{ii}|$, and $C$ is from (34);
  2. 2.
    If $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds, then the cumulative privacy budget remains finite as $T\rightarrow\infty$.
@@
$\quad+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}.$
Therefore, it can be obtained from (34) and (36) that the sensitivity $\Delta^{k}$ is bounded by
$\Delta^{k+1}$ $\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}$ (37)
$\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+C\gamma^{k}\chi^{k}.$
According to Lemma 2 and (37), the first statement can be obtained.
Lemma 3 is exploited to prove the second statement of Theorem 3. Specifically, based on (37) and the properties of $\chi^{k}$ and $\gamma^{k}$, Lemma 3 can be used to conclude that there exists some constant $\bar{C}$ such that the sensitivity $\Delta^{k}$ satisfies $\Delta^{k}\leq\bar{C}\gamma^{k}$. It can be further obtained from Lemma 2 that $\epsilon\leq\sum_{k=1}^{T}\frac{\bar{C}\gamma^{k}}{\nu^{k}}$. Thus, if $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds (i.e., the sequence $\{\frac{\gamma^{k}}{\nu^{k}}\}$ is summable), then $\epsilon$ will be finite even when $T\rightarrow\infty$. ∎

### v3

@@ III-B Convergence Analysis
Based on Lemma 1 in [NedicTAC2010] and the update law of $\lambda_{i}^{k}$ in (16), it can be obtained that for any $\lambda\in\mathbb{R}_{+}^{Np}$, $\sum_{i=1}^{M}\|\lambda_{i}^{k\!+\!1}\!-\!\lambda\|^{2}=\sum_{i=1}^{M}\|\Pi_{\mathbb{R}^{Np}_{+}}[\lambda_{i}^{k}\!+\!\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}\!-\!\lambda_{i}^{k})\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})]\!-\!\lambda\|^{2}\leq\sum_{i=1}^{M}\|\lambda_{i}^{k}\!+\!\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}\!-\!\lambda_{i}^{k})\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\!-\!\lambda\|^{2}\leq\sum_{i=1}^{M}\|\lambda_{i}^{k}\!+\!\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\lambda_{j}^{k}\!+\!\zeta_{j}^{k}\!-\!\lambda_{i}^{k})\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\!-\!\lambda\|^{2}\leq\sum_{i=1}^{M}\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}\!-\!\lambda\!+\!\chi^{k}\xi_{i}^{k}\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\|^{2}$, where $w_{ij}^{k}$ and $\xi_{i}^{k}$ are defined as
$w_{ii}^{k}=1+\chi^{k}L_{ii},\qquad w_{ij}^{k}=\chi^{k}L_{ij},\qquad\xi_{i}^{k}=\sum_{j\in\mathcal{N}_{i}}L_{ij}\zeta_{j}^{k}.$ (19)
Using $w_{ij}^{k}\lambda_{j}^{k}\!-\!\lambda=(\bar{\lambda}^{k}\!-\!\lambda)+(w_{ij}^{k}\lambda_{j}^{k}\!-\!\bar{\lambda}^{k})$, it can be further concluded that
$\sum_{i=1}^{M}\!\|\lambda_{i}^{k\!+\!1}\!-\!\lambda\|^{2}\!\!\leq\!\sum_{i=1}^{M}\!\left(\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\lambda_{j}^{k}\!-\!\lambda\|^{2}\!\!\!+\!\|\chi^{k}\xi_{i}^{k}\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\|^{2}\right$. (20)
$\hskip 18.49988pt\left.+2\!\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\lambda_{j}^{k}\!-\!\lambda\right)^{\top}\!\left(\chi^{k}\xi_{i}^{k}\right)\right.\left.\!\!+\!2\left(\bar{\lambda}^{k}\!-\!\lambda\right)^{\top}\!\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\right)\right$.
$\hskip 18.49988pt\left.+2\!\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\lambda_{j}^{k}\!-\!\bar{\lambda}^{k}\right)^{\top}\!\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k\!+\!1})\right)\right).$
@@
By (21) and by the convexity of $\|\cdot\|^{2}$, we have that
$\sum_{i=1}^{M}\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\lambda_{j}^{k}\!-\!\lambda\|^{2}\!=\!\sum_{i=1}^{M}\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\left(\lambda_{j}^{k}\!-\!\lambda\right)\|^{2}$ (23)
$\leq\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}\cup\{i\}}\!\!\!w_{ij}^{k}\|\left(\lambda_{j}^{k}-\lambda\right)\|^{2}\leq\sum_{i=1}^{M}\|\lambda_{i}^{k}-\lambda\|^{2}.$
@@
$\hskip 18.49988pt\left.+J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})\right)$
@@
$\sum_{i=1}^{M}\mathbb{E}\left[\|\lambda_{i}^{k+1}\!-\!\lambda\|^{2}|\mathcal{F}^{k}\right]$ (25)
$\leq$ $\sum_{i=1}^{M}\|\lambda_{i}^{k}\!-\!\lambda\|^{2}\!+\!d^{k}\!+\!2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\bar{\lambda}^{k})\!-\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k\!+\!1}\},\lambda)\right),$
where $d^{k}\!=\!(\chi^{k})^{2}\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}+M(\gamma^{k})^{2}C_{g}^{2}+6C_{g}\gamma^{k}\sum_{i=1}^{M}\|\lambda_{i}^{k}-\bar{\lambda}^{k}\|$. Based on Assumption 4, Theorem 1, and the conditions for $\chi^{k}$ and $\gamma^{k}$, it can be concluded that $d^{k}$ is summable, i.e., $\sum_{k=0}^{\infty}d^{k}<\infty$.
Plugging the optimal primal-dual pair $(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ into (25) and utilizing the Saddle-Point Theorem (10), we can arrive at
$\sum_{i=1}^{M}\mathbb{E}\left[\|\lambda_{i}^{k+1}-\lambda^{*}\|^{2}|\mathcal{F}^{k}\right]$ (26)
$\leq$ $\sum_{i=1}^{M}\|\lambda_{i}^{k}\!-\!\lambda^{*}\|^{2}\!\!+\!d^{k}\!+\!2\gamma^{k}\left(\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})\!-\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\!\right),$
$\sum_{i=1}^{M}\mathbb{E}\left[\|\lambda_{i}^{k+1}-\lambda^{*}\|^{2}|\mathcal{F}^{k}\right]$ (27)
$\leq$ $\sum_{i=1}^{M}\!\|\lambda_{i}^{k}\!-\!\lambda^{*}\|^{2}\!\!+\!d^{k}\!+\!2\gamma^{k}\!\left(\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\!-\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k\!+\!1}\},\lambda^{*})\!\right).$
According to Lemma 1 and $\sum_{k=0}^{\infty}d^{k}<\infty$, it can be concluded that $\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right)$ in (26) and $\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right)$ in (27) satisfy the conditions for $\phi^{k}$ in Lemma 1, i.e., the following relationships hold almost surely:
$\sum_{k=1}^{\infty}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right)<\infty,$ (28)
$\sum_{k=1}^{\infty}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right)<\infty.$
Since $\gamma^{k}$ is non-summable, we have that $\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and $\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})$ converge to zero almost surely. ∎
@@
The weakening factor sequence $\{\chi^{k}\}$ is employed to mitigate the influence of persistent DP noise, and the step-size sequence $\{\gamma^{k}\}$ should be appropriately selected to work with $\{\chi^{k}\}$ for guaranteed convergence. The conditions for the sequences $\{\chi^{k}\}$ and $\{\gamma^{k}\}$ in Theorems 1 and 2 can be satisfied, e.g., by selecting $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\gamma^{k}=\frac{c_{4}}{1+c_{5}k}$ with any $c_{1}>0$, $c_{2}>0$, $0.5<c_{3}<1$, $c_{4}>0$, and $c_{5}>0$. Note that the design of $\chi^{k}$ in this example is identical to the one in Remark 1. Therefore, the sequences $\{\chi^{k}\}$, $\{\gamma^{k}\}$, and $\{\nu^{k}\}$ can be meticulously tailored to meet all conditions required by Assumption 4 and Theorems 1, 2.
@@ III-C Privacy Analysis
Using the adjacency concept defined in Definition 1, we establish two adjacent distributed optimization problems, denoted as $\mathcal{P}$ and $\mathcal{P}^{\prime}$. There is only one signal that differs between these two problems, and without loss of generality we denote it as $g_{i}$ in $\mathcal{P}$ and $g^{\prime}_{i}$ in $\mathcal{P}^{\prime}$. Per the third condition of Definition 1, the signals $g_{i}$ and $g^{\prime}_{i}$ should exhibit similar behavior near the optimal solution. Specifically, $g_{i}$ and $g^{\prime}_{i}$ should converge toward each other if the algorithm guarantees convergence to the optimal solution. Leveraging the proven convergence from Theorem 2, we formalize this condition by requiring the existence of a constant $C>0$ such that:
$\|g_{i}(\tilde{\bm{u}}^{k+1}_{i})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}\leq C\chi^{k}$ (29)
@@
For Algorithm 3, an execution is represented as $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\ldots\}$ with $\vartheta^{k}=\lambda^{k}=\begin{bmatrix}(\lambda_{1}^{k})^{\top},\cdots,(\lambda_{M}^{k})^{\top}\end{bmatrix}^{\top}$. An observation sequence is denoted as $\mathcal{O}=\{o^{0},o^{1},\ldots\}$ with $o^{k}=\hat{\lambda}^{k}=\begin{bmatrix}(\hat{\lambda}_{1}^{k})^{\top},\cdots,(\hat{\lambda}_{M}^{k})^{\top}\end{bmatrix}^{\top}$ (note that $\hat{\lambda}_{i}^{k}=\lambda_{i}^{k}+\zeta_{i}^{k}$, as detailed in Algorithm 3). Similar to the sensitivity metric for constraint-free distributed optimization in [Huang2015], we formulate the sensitivity of Algorithm 3 as follows:
@@
$\Delta^{k}=\sup\limits_{\mathcal{O}\in\mathbb{O}}\left\{\sup\limits_{\vartheta\in\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O}),\>\vartheta^{{}^{\prime}}\in\mathcal{R}_{\mathcal{P}^{{}^{\prime}},\vartheta^{0}}^{-1}(\mathcal{O})}\hskip-8.5359pt\|\vartheta^{k}-\vartheta^{\prime k}\|_{1}\right\},$ (30)
@@
In Algorithm 3, at each iteration $k$, if each subsystem’s DP noise vector $\zeta_{i}^{k}\in\mathbb{R}^{Np}$ comprises NpNp independent Laplace noises with parameter $\nu^{k}$, satisfying $\sum_{k=1}^{T_{0}}\frac{\Delta^{k}}{\nu^{k}}\leq\bar{\epsilon}$ for some $\bar{\epsilon}>0$, then Algorithm 3 achieves $\epsilon$-differential privacy with the cumulative privacy level for iterations $0\leq k\leq T_{0}$ less than $\bar{\epsilon}$.
@@
The proof of this lemma follows the same reasoning as that of Lemma 2 in [Huang2015]. ∎
@@
(Lemma 4, [WangTAC2023]) Let $\{\psi^{k}\}$ be a non-negative sequence, and $\{a^{k}\}$ and $\{\varpi^{k}\}$ be positive sequences satisfying $\sum_{k=0}^{\infty}a^{k}=\infty$, $\lim_{k\rightarrow\infty}a^{k}=0$, and $\frac{\varpi^{k}}{a^{k}}$ converges to zero with a polynomial rate. If there exists a $\bar{K}\geq 0$ such that $\psi^{k+1}\leq(1-a^{k})\psi^{k}+\varpi^{k}$ holds for all $k\geq\bar{K}$, then it follows that $\psi^{k}\leq\bar{C}\frac{\varpi^{k}}{a^{k}}$ for all $k$, with $\bar{C}$ being some constant.
@@
Suppose the conditions of Theorem 1 hold. If every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$ satisfies Assumption 4, then the following results hold: $\left.1\right)$ For any finite number of iterations $T$, Algorithm 3 ensures $\epsilon$-differential privacy, and the cumulative privacy budget is bounded by $\epsilon\leq\sum_{k=1}^{T}\frac{C\nu^{k}}{\nu^{k}}$, where $\varsigma^{k}:=\sum_{s=1}^{k-1}\Pi_{q=s}^{k-1}(1\!-\!\chi^{q}\bar{L})\gamma^{s-1}\chi^{s-1}\!+\!\gamma^{k-1}\chi^{k-1}$, $\bar{L}:=\min_{i\in\mathbb{Z}_{1}^{M}}|L_{ii}|$, and $C$ is from (29); $\left.2\right)$ If $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds, the cumulative privacy budget remains finite as $T\rightarrow\infty$.
@@
$\qquad+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}.$
From (29) and (31), it follows that
$\Delta^{k+1}\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+C\gamma^{k}\chi^{k}.$ (32)
Using Lemma 2 and (32), the first statement is established.
Lemma 3 is applied to prove the second statement of Theorem 3. Specifically, based on (32) and the properties of $\chi^{k}$ and $\gamma^{k}$, Lemma 3 implies that there exists some constant $\bar{C}$ such that the sensitivity $\Delta^{k}$ satisfies $\Delta^{k}\leq\bar{C}\gamma^{k}$. It can be further obtained from Lemma 2 that $\epsilon\leq\sum_{k=1}^{T}\frac{\bar{C}\gamma^{k}}{\nu^{k}}$. Thus, if $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds (i.e., the sequence $\{\frac{\gamma^{k}}{\nu^{k}}\}$ is summable), then $\epsilon$ will be finite even when $T\rightarrow\infty$. ∎
For $\epsilon$-differential privacy (see Definition 2), a smaller $\epsilon$ indicates a better extent of privacy preservation. According to Theorem 3, for given $C$ and $\varsigma^{k}$, a higher noise level $\nu^{k}$ results in a smaller $\epsilon$, thereby enhancing privacy protection.
Remark 3.
The DP noise injection mechanism in Algorithm 3 is computationally efficient and easy to implement. To mitigate the impact of DP noise, careful design of the weakening factor and step-size sequences is essential. When the DP noise intensity is high, the algorithm may require more iterations to converge compared to non-privacy-preserving methods. This trade-off is necessary to achieve both $\epsilon$-differential privacy and provable convergence to the optimal solution.